Brooks Sporting Inc. is prepared to report the following 2019 income statement (shown in thousands of dollars).

Sales $10,200 Operating costs including depreciation 7,344 EBIT $2,856 Interest 396 EBT $2,460 Taxes (25%) 615 Net income $1,845

Prior to reporting this income statement, the company wants to determine its annual dividend.

A.)The company has 500,000 shares of common stock outstanding, and its stock trades at $40 per share. The company had a 35% dividend payout ratio in 2018. If Brooks wants to maintain this payout ratio in 2019, what will be its per-share dividend in 2019? Do not round intermediate calculations. Round your answer to the nearest cent. $

B.)If the company maintains this 35% payout ratio, what will be the current dividend yield on the company's stock? Do not round intermediate calculations. Round your answer to two decimal places. %

C.) The company reported net income of $1.65 million in 2018. Assume that the number of shares outstanding has remained constant. What was the company's per-share dividend in 2018? Do not round intermediate calculations. Round your answer to the nearest cent. $

D.) As an alternative to maintaining the same dividend payout ratio, Brooks is considering maintaining the same per-share dividend in 2019 that it paid in 2018. If it chooses this policy, what will be the company's dividend payout ratio in 2019? Do not round intermediate calculations. Round your answer to two decimal places. %
Answer #1
a.Per share dividend in 2019 = 1845*35%/500 1.2915
b.Dividend yield = Expected Dividend/Current Price 3.2288%
c.Per share dividend in 2018 = 1,650,000*35%/500,000 1.155
d.Payout ratio 31.3008%

=1845*35%/500 a.Per share dividend in 2019 = 1845*35%/500 2 3 b.Dividend yield = Expected Dividend/Current Price =B1/40 %3D 5